# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ 5.3 Tractable Subcases
5.4 Bridge from Stochastic
@@
Proposition 5.14 (Stochastic Sufficiency Does Not Imply Sequential Sufficiency).
There exist instances where $I$ is sufficient in the stochastic one-shot sense but fails to be sufficient once temporal dynamics are introduced.
(L: DQ62.)
@@
Construct a process in which the same one-step expected optimizer is preserved under $I$, but hidden-state memory affects future information states and therefore optimal policies. The temporal dependence blocks one-shot sufficiency transfer.  $\blacksquare$
@@ 6.1 Complexity Matrix
Stochastic decisiveness P (explicit), PP-hard (succinct) PP-hard, in $\textsf{NP}^{\textsf{PP}}$ PP-hard, in $\textsf{NP}^{\textsf{PP}}$
@@
This table is the paper’s compact summary. The stochastic regime contains two parallel predicates with distinct complexity profiles: preservation and decisiveness. In the static regime, sufficiency and minimum sufficiency are coNP-complete and anchor sufficiency is $\Sigma_{2}^{P}$-complete. In the stochastic regime, preservation (stochastic sufficiency) has explicit-state tractability and bridge theorems back to static sufficiency and the optimizer quotient; the minimum and anchor preservation variants are also polynomial-time in explicit state, inherit coNP-completeness and $\Sigma_{2}^{P}$-completeness under full support, and admit support-sensitive obstruction and bridge lemmas beyond full support, while the full general-distribution classification remains open. Decisiveness (stochastic decisiveness) has the sharper succinct-encoding classification: polynomial-time in the explicit-state model, PP-hardness in the succinct model, and paper-level $\textsf{NP}^{\textsf{PP}}$ membership for the anchor and minimum variants. The artifact mechanizes the finite witness/checking schemas and reduction cores underlying those upper bounds. In the sequential regime, sufficiency, minimum, and anchor queries are PSPACE-complete.
Preservation Boundary.
For stochastic sufficiency in the preservation sense, this paper proves explicit-state tractability for the preservation family (Theorem 4.5 and Theorem 4.13), static-to-stochastic bridge theorems (Propositions 4.6 and 4.7), quotient equivalence under full support (Proposition 4.8), full-support inheritance for the minimum and anchor variants (Theorem 4.11), and support-sensitive obstruction/bridge lemmas beyond full support (Proposition 4.14). Outside the full-support setting, this paper does not claim a succinct-encoding hardness classification for the preservation predicate or a full general-distribution classification for its minimum and anchor variants. The PP-hardness results in the stochastic row concern decisiveness and its variants, not preservation itself.
The decisiveness results combine standard hardness claims with mechanized upper-bound packages. The artifact verifies the finite witness/checking schemata and reduction cores used by the stochastic oracle-class arguments, while the corresponding oracle-class memberships are proved in the paper text by standard complexity-theoretic reasoning.
The stochastic anchor and minimum decisiveness queries are not presently collapsed to PP-completeness: Theorem 4.24 places them in $\textsf{NP}^{\textsf{PP}}$, while Theorems 4.21 and 4.22 give PP-hardness. The artifact verifies the finite witness/checking structure and explicit-state bounded-witness recoveries behind that upper-bound story, and it also packages existential-majority hardness/completeness surrogates for the anchor and decisiveness query families (L: EH2-10.) The oracle-class memberships themselves remain paper-level proofs under the succinct encoding. By contrast, the sequential anchor and minimum variants admit paper-level PSPACE membership via guess-and-verify over anchors or coordinate subsets together with the sequential sufficiency verifier.
Theorem 6.1 (Regime-Sensitive Complexity Matrix).
Under the encoding conventions of Section 2.4, the classification is as follows. In the static regime, sufficiency and minimum sufficiency are coNP-complete and anchor sufficiency is $\Sigma_{2}^{P}$-complete. In the stochastic regime, preservation (stochastic sufficiency) has a polynomial-time base query under explicit-state encoding with proved bridges back to static sufficiency and the optimizer quotient; the minimum and anchor preservation variants are also polynomial-time under explicit-state encoding, are coNP-complete and $\Sigma_{2}^{P}$-complete under full support, and satisfy the support-sensitive partial results of Proposition 4.14. Decisiveness (stochastic decisiveness) has a polynomial-time base query under explicit-state encoding and is PP-hard under the succinct encoding; the anchor and minimum decisiveness queries are PP-hard and lie in $\textsf{NP}^{\textsf{PP}}$, with those oracle-class memberships proved in the paper text and their finite witness/checking cores mechanized. In the sequential regime, sufficiency, minimum, and anchor queries are PSPACE-complete.
(L: DC100, DQ1, DQ17-19, OU1-2, OU11-12, EH4-10.)
@@
The matrix entries are exactly Theorems 3.4, 3.5, 3.6, 4.5, 4.11, 4.13, Proposition 4.14, 4.19, 4.21, 4.22, 4.24, 5.8, 5.10, and 5.11, together with the stochastic sufficiency bridge package. The strict regime comparisons are isolated separately in Theorems 6.2 and 6.3.  $\blacksquare$
6.2 Strict Regime Gaps
Theorem 6.2 (Static-to-Stochastic Complexity Gap).
Assuming $\textsf{coNP}\neq\textsf{PP}$, the succinct stochastic decisiveness classification is strictly harder in worst-case complexity than the static sufficiency classification.
(L: DQ18.)
@@ 6.2 Strict Regime Gaps
Static sufficiency checking is coNP-complete (Theorem 3.4), while stochastic decisiveness is PP-hard under the succinct encoding (Theorem 4.19). Under the assumption $\textsf{coNP}\neq\textsf{PP}$, this yields a strict complexity gap between the two classifications.  $\blacksquare$
Theorem 6.3 (Stochastic-to-Sequential Complexity Gap).
Assuming $\textsf{PP}\neq\textsf{PSPACE}$, the sequential sufficiency classification is strictly harder in worst-case complexity than the stochastic decisiveness classification.
(L: DQ19.)
@@
Stochastic decisiveness is PP-hard under the succinct encoding (Theorem 4.19), while sequential sufficiency is PSPACE-complete (Theorem 5.8). The class gap yields a strict complexity separation between the two classifications under the assumption $\textsf{PP}\neq\textsf{PSPACE}$.  $\blacksquare$
6.3 Integrity-Resource Bounds
The regime matrix has a methodological consequence as well: as one moves upward in the hierarchy, exact certification becomes less compatible with fixed polynomial resource budgets.
Proposition 6.4 (Abstention Frontier Expands with Regime Complexity).
$\text{static}\to\text{stochastic}\to\text{sequential}.$
(L: IC32, DQ2.)
Proof.
As the decision predicate moves from coNP to PP to PSPACE hardness, exact worst-case certification requires strictly stronger resources. Fixed polynomial resources therefore force abstention on a larger instance family in higher regimes.  $\blacksquare$
@@ 6.3 Integrity-Resource Bounds
7 Encoding Dichotomy and ETH Lower Bounds
The regime hierarchy identifies where exact relevance certification sits in the standard complexity landscape. This section extracts the complementary algorithmic message: under the encodings of Section 2.4, there is a strong contrast between explicit-state instances whose true relevant support is small and succinct instances whose relevant support is extensive. The former admit direct algorithms; the latter inherit exponential lower bounds under ETH. The practical takeaway is cross-model rather than single-model: logarithmic relevant support can still be certified exactly in polynomial time under explicit encodings, while linear relevant support under succinct encodings already forces ETH-conditioned exponential cost.
@@ 7 Encoding Dichotomy and ETH Lower Bounds
Part 1 is an explicit-state upper bound (time polynomial in $|S|$). Part 2 is a succinct-encoding lower bound under ETH (time exponential in $n$). The encodings are defined in Section 2.4.
@@
Explicit-state $k^{*}=O(\log N)$ exact certification is polynomial in $N$ by projection enumeration Theorem 7.1(1)
Succinct $k^{*}=\Omega(n)$ exact certification inherits an ETH-conditioned $2^{\Omega(n)}$ lower bound Theorem 7.1(2)
Boundary summary. The tractable-hard contrast is controlled jointly by encoding and by the size of the true sufficient support.
Theorem 7.1 (Encoding-Sensitive Contrast).
@@
    Logarithmic case (explicit-state upper bound): If $k^{*}=O(\log N)$, then SUFFICIENCY-CHECK is solvable in polynomial time in $N$ under the explicit-state encoding.
@@
7.1 The ETH Reduction Chain
The lower bound in Part 2 of Theorem 7.1 follows from a short standard ETH transfer. Under ETH, 3-SAT requires $2^{\Omega(n)}$ time. Negation gives a linear-time reduction from 3-SAT to TAUTOLOGY. The static reduction of Theorem 3.4 then maps an $n$-variable tautology instance to a sufficiency instance with the same asymptotic coordinate count. Therefore, if Sufficiency-Check were solvable in $2^{o(n)}$ time under the succinct encoding, then so would TAUTOLOGY and hence 3-SAT, contradicting ETH.
Definition 7.2 (Exponential Time Hypothesis (ETH)).
@@ 7.1 The ETH Reduction Chain
Proposition 7.3 (Tight Constant).
The reduction in Theorem 3.4 preserves the number of variables up to an additive constant: an $n$-variable formula yields an $(n+1)$-coordinate decision problem. Therefore, the constant $c$ in the $2^{n/c}$ lower bound is asymptotically 1:
@@
The TAUTOLOGY reduction (Theorem 3.4) constructs:
@@
7.2 Cross-Model Contrast
Corollary 7.4 (Cross-Model Contrast for Exact Certification).
Across the explicit-state and succinct encodings of Section 2.4, exact certification exhibits the following contrast:
  • If $k^{*}=O(\log N)$, SUFFICIENCY-CHECK is polynomial in $N$ under the explicit-state encoding
@@ 7.2 Cross-Model Contrast
The logarithmic case (Part 1 of Theorem 7.1) gives polynomial time when $k^{*}=O(\log N)$. More precisely, when $k^{*}\leq c\log N$ for constant $c$, the algorithm runs in time $O(N^{c}\cdot\text{poly}(n))$.
@@
The transition point inside the explicit-state bound is where $2^{k^{*}}=N^{k^{*}/\log N}$ stops being polynomial in $N$, i.e., when $k^{*}=\omega(\log N)$. For Boolean coordinate spaces, however, the present theorem should be read only as a cross-encoding contrast, because $\log N=n$ and the succinct lower bound is parameterized directly by the succinct variable count.  $\blacksquare$
Remark 7.5 (Asymptotic Tightness Within Each Encoding).
Under ETH, the lower bound is asymptotically tight in the succinct encoding. The explicit-state upper bound is tight in the sense that it matches enumeration complexity in $N$.
@@
Remark 7.6 (Circuit Model Formalization).
The ETH lower bound is stated in the succinct circuit encoding of Section 2.4, where the utility function $U:A\times S\to\mathbb{R}$ is represented by a Boolean circuit computing $\mathbf{1}[U(a,s)>\theta]$ for threshold comparisons. In this model:
@@
8 Tractable Special Cases
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly.
@@ 8 Tractable Special Cases
Restriction Why it helps Complexity
@@
Tractability map. Each nontrivial restriction above removes a structural hardness source. Additional simple cases include bounded actions ($O(|S|^{2}\cdot k^{2})$ when $|A|$ is fixed, separable utility ($O(1)$), and constant or dominated optimal sets ($O(1)$). Formal definitions for all cases are retained in the subsections below.
Theorem 8.1 (Tractable Subcases).
@@
    Low tensor rank: The utility admits a rank-$R$ decomposition $U(a,s)=\sum_{r=1}^{R}w_{r}\cdot f_{r}(a)\cdot\prod_{i}g_{ri}(s_{i})$. The artifact certifies reduction to factored computation with bound $O(|A|\cdot R\cdot n)$. (L: DQ77.)
@@
    Tree-structured dependencies: Coordinates are ordered such that $s_{i}$ depends only on $(s_{1},\ldots,s_{i-1})$. The artifact certifies an explicit sufficiency checker for this class. (L: DQ63.)
@@
    Bounded treewidth: The interaction graph has treewidth $\leq w$ and utility factors over edges. The artifact certifies reduction to bounded-treewidth CSP together with bound $O(n\cdot k^{w+1})$ inherited from standard CSP algorithm. (L: DQ81.)
@@
    Coordinate symmetry: Utility is invariant under coordinate permutations. The artifact certifies reduction to orbit-type representatives together with resulting orbit-count bound $O\bigl(\binom{d+k-1}{k-1}^{2}\bigr)$. (L: DQ85.)
@@ 8.1 Bounded Actions
Definition 8.2 (Bounded Action Problem).
@@
Theorem 8.3 (Bounded Actions Tractability).
@@
    Compute $\text{Opt}(s)$ for all $s\in S$ in $O(|S|\cdot k)$ time.
@@
    For each pair $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$, compare $\text{Opt}(s)$ and $\text{Opt}(s^{\prime})$ in $O(k^{2})$ time.
@@
8.2 Separable Utility (Rank 1)
@@
Definition 8.4 (Separable Utility).
@@
Theorem 8.5 (Separable Tractability).
@@
$\text{Opt}(s)=\arg\max_{a\in A}[f(a)+g(s)]=\arg\max_{a\in A}f(a)$
The optimal action depends only on $f$, not on $s$. Hence $\text{Opt}(s)=\text{Opt}(s^{\prime})$ for all $s,s^{\prime}$, making any $I$ (including $\emptyset$) sufficient.  $\blacksquare$
8.3 Low Tensor Rank

### 31 Mar 2026 v2

@@ V-C Tractable Subcases
V-D Bridge from Stochastic
@@
Proposition V.14footnotemark: footnotetext: : DQ62 (Stochastic Sufficiency Does Not Imply Sequential Sufficiency).
There exist instances where II is sufficient in the stochastic one-shot sense but fails to be sufficient once temporal dynamics are introduced.
@@
Construct a process in which the same one-step expected optimizer is preserved under II, but hidden-state memory affects future information states and therefore optimal policies. The temporal dependence blocks one-shot sufficiency transfer.  $\blacksquare$
@@ VI-A Complexity Matrix
Stochastic decisiveness P (explicit), $P$-hard (succinct) $P$-hard, in $\textsf{NP}^{\textsf{PP}}$ $P$-hard, in $\textsf{NP}^{\textsf{PP}}$
@@
TABLE III: Complexity matrix across regimes and query types.
Table III is the main statement-level summary. The stochastic row is intentionally split: preservation and decisiveness are distinct predicates with different complexity behavior. Open cells are explicit scope boundaries.
Theorem VI.1footnotemark: footnotetext: : DC100, DQ1, DQ17-19, OU1-2, OU11-12, EH4-10 (Regime-Sensitive Complexity Matrix).
Under the encoding conventions of Section II-D, the classification is as follows. In the static regime, sufficiency and minimum sufficiency are coNP-complete and anchor sufficiency is $\Sigma_{2}^{P}$-complete. In the stochastic regime, preservation (stochastic sufficiency) has a polynomial-time base query under explicit-state encoding with proved bridges back to static sufficiency; the minimum and anchor preservation variants are also polynomial-time under explicit-state encoding, are coNP-complete and $\Sigma_{2}^{P}$-complete under full support, and satisfy the support-sensitive partial results of Proposition IV.16. Decisiveness (stochastic decisiveness) has a polynomial-time base query under explicit-state encoding and is $P$-hard under the succinct encoding; the anchor and minimum decisiveness queries are $P$-hard and lie in $\textsf{NP}^{\textsf{PP}}$, with those oracle-class memberships proved in the paper text and their finite witness/checking cores mechanized. In the sequential regime, sufficiency, minimum, and anchor queries are PSPACE-complete.
@@
Immediate by table lookup from Theorems III.4, III.5, III.6, IV.7, IV.13, IV.15, Proposition IV.16, IV.21, IV.23, IV.24, IV.26, V.8, V.10, and V.11.  $\blacksquare$
VI-B Strict Regime Gaps
Theorem VI.2footnotemark: footnotetext: : DQ18 (Static-to-Stochastic Complexity Gap).
Assuming $\mathrm{coNP}\neq\mathrm{PP}$, the succinct stochastic decisiveness classification is strictly harder in worst-case complexity than the static sufficiency classification.
@@ VI-B Strict Regime Gaps
Static sufficiency checking is coNP-complete (Theorem III.4), while stochastic decisiveness is $P$-hard under the succinct encoding (Theorem IV.21). Under the assumption $\mathrm{coNP}\neq\mathrm{PP}$, this yields a strict complexity gap between the two classifications.  $\blacksquare$
Theorem VI.3footnotemark: footnotetext: : DQ19 (Stochastic-to-Sequential Complexity Gap).
Assuming $\mathrm{PP}\neq\mathrm{PSPACE}$, the sequential sufficiency classification is strictly harder in worst-case complexity than the stochastic decisiveness classification.
@@
Stochastic decisiveness is $P$-hard under the succinct encoding (Theorem IV.21), while sequential sufficiency is PSPACE-complete (Theorem V.8). The class gap yields a strict complexity separation between the two classifications under the assumption $\mathrm{PP}\neq\mathrm{PSPACE}$.  $\blacksquare$
@@
VII Encoding Dichotomy and ETH Lower Bounds
The regime hierarchy identifies where exact relevance certification sits in the standard complexity landscape. This section extracts the complementary algorithmic message: under the encodings of Section II-D, there is a strong contrast between explicit-state instances whose true relevant support is small and succinct instances whose relevant support is extensive. The former admit direct algorithms; the latter inherit exponential lower bounds under ETH. The practical takeaway is cross-model rather than single-model: logarithmic relevant support can still be certified exactly in polynomial time under explicit encodings, while linear relevant support under succinct encodings already forces ETH-conditioned exponential cost.
@@ VII Encoding Dichotomy and ETH Lower Bounds
Part 1 is an explicit-state upper bound (time polynomial in $|S|$). Part 2 is a succinct-encoding lower bound under ETH (time exponential in $n$). The encodings are defined in Section II-D.
@@
Explicit-state $k^{*}=O(\log N)$ exact certification is polynomial in NN by projection enumeration Theorem VII.1(1)
Succinct $k^{*}=\Omega(n)$ exact certification inherits an ETH-conditioned $2^{\Omega(n)}$ lower bound Theorem VII.1(2)
TABLE IV: Boundary summary. The tractable-hard contrast is controlled jointly by encoding and by the size of the true sufficient support.
Theorem VII.1footnotemark: footnotetext: : DQ26, DQ29-30 (Encoding-Sensitive Contrast).
@@
    Logarithmic case (explicit-state upper bound): If $k^{*}=O(\log N)$, then SUFFICIENCY-CHECK is solvable in polynomial time in NN under the explicit-state encoding.
@@
VII-A The ETH Reduction Chain
The lower bound in Part 2 of Theorem VII.1 follows from a short standard ETH transfer. Under ETH, 3-SAT requires $2^{\Omega(n)}$ time. Negation gives a linear-time reduction from 3-SAT to TAUTOLOGY. The static reduction of Theorem III.4 then maps an $n$-variable tautology instance to a sufficiency instance with the same asymptotic coordinate count. Therefore, if Sufficiency-Check were solvable in $2^{o(n)}$ time under the succinct encoding, then so would TAUTOLOGY and hence 3-SAT, contradicting ETH.
Definition VII.2 (Exponential Time Hypothesis (ETH)).
@@ VII-A The ETH Reduction Chain
Proposition VII.3footnotemark: footnotetext: : DQ27, DQ30 (Tight Constant).
The reduction in Theorem III.4 preserves the number of variables up to an additive constant: an $n$-variable formula yields an $(n+1)$-coordinate decision problem. Therefore, the constant $c$ in the $2^{n/c}$ lower bound is asymptotically 1:
@@
The TAUTOLOGY reduction (Theorem III.4) constructs:
@@
VII-B Cross-Model Contrast
Corollary VII.4footnotemark: footnotetext: : DQ3, DQ26 (Cross-Model Contrast for Exact Certification).
Across the explicit-state and succinct encodings of Section II-D, exact certification exhibits the following contrast:
  • If $k^{*}=O(\log N)$, SUFFICIENCY-CHECK is polynomial in NN under the explicit-state encoding
@@ VII-B Cross-Model Contrast
The logarithmic case (Part 1 of Theorem VII.1) gives polynomial time when $k^{*}=O(\log N)$. More precisely, when $k^{*}\leq c\log N$ for constant $c$, the algorithm runs in time $O(N^{c}\cdot\text{poly}(n))$.
@@
The transition point inside the explicit-state bound is where $2^{k^{*}}=N^{k^{*}/\log N}$ stops being polynomial in NN, i.e., when $k^{*}=\omega(\log N)$. For Boolean coordinate spaces, however, the present theorem should be read only as a cross-encoding contrast, because $\log N=n$ and the succinct lower bound is parameterized directly by the succinct variable count.  $\blacksquare$
Remark VII.5 (Asymptotic Tightness Within Each Encoding).
Under ETH, the lower bound is asymptotically tight in the succinct encoding. The explicit-state upper bound is tight in the sense that it matches enumeration complexity in NN.
@@
Remark VII.6 (Circuit Model Formalization).
The ETH lower bound is stated in the succinct circuit encoding of Section II-D, where the utility function $U:A\times S\to\mathbb{R}$ is represented by a Boolean circuit computing $\mathbf{1}[U(a,s)>\theta]$ for threshold comparisons. In this model:
@@
VIII Tractable Special Cases
We distinguish the encodings of Section II-D. The tractability results below state the model assumption explicitly.
@@ VIII Tractable Special Cases
Restriction Algorithmic mechanism Complexity
Bounded actions brute-force enumeration is polynomial in $|S|$ when $|A|$ constant $O(|S|^{2}\cdot k^{2})$
Separable utility optimal action independent of state $O(1)$
@@
TABLE V: Tractability map. Each restriction above removes a structural hardness source. Formal definitions for all cases are retained in the subsections below.
Theorem VIII.1 (Tractable Subcases).
@@
    Low tensor rank: The utility admits a rank-$R$ decomposition $U(a,s)=\sum_{r=1}^{R}w_{r}\cdot f_{r}(a)\cdot\prod_{i}g_{ri}(s_{i})$. The artifact certifies reduction to factored computation with bound $O(|A|\cdot R\cdot n)$. 535353: DQ77.
@@
    Tree-structured dependencies: Coordinates are ordered such that $s_{i}$ depends only on $(s_{1},\ldots,s_{i-1})$. The artifact certifies an explicit sufficiency checker for this class. 545454: DQ63.
@@
    Bounded treewidth: The interaction graph has treewidth $\leq w$ and utility factors over edges. The artifact certifies reduction to bounded-treewidth CSP together with bound $O(n\cdot k^{w+1})$ inherited from standard CSP algorithm. 555555: DQ81.
@@
    Coordinate symmetry: Utility is invariant under coordinate permutations. The artifact certifies reduction to orbit-type representatives together with resulting orbit-count bound $O\bigl(\binom{d+k-1}{k-1}^{2}\bigr)$. 565656: DQ85.
@@ VIII-A Bounded Actions
Definition VIII.2 (Bounded Action Problem).
@@
Theorem VIII.3footnotemark: footnotetext: : DQ51, DQ72 (Bounded Actions Tractability).
@@
    Compute $\mathrm{Opt}(s)$ for all $s\in S$ in $O(|S|\cdot k)$ time.
@@
    For each pair $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$, compare $\mathrm{Opt}(s)$ and $\mathrm{Opt}(s^{\prime})$ in $O(k^{2})$ time.
@@
VIII-B Separable Utility (Rank 1)
@@
Definition VIII.4 (Separable Utility).
@@
Theorem VIII.5footnotemark: footnotetext: : DQ57, DQ74 (Separable Tractability).
@@
$\mathrm{Opt}(s)=\arg\max_{a\in A}[f(a)+g(s)]=\arg\max_{a\in A}f(a)$
The optimal action depends only on $f$, not on $s$. Hence $\mathrm{Opt}(s)=\mathrm{Opt}(s^{\prime})$ for all $s,s^{\prime}$, making any II (including $\emptyset$) sufficient.  $\blacksquare$
VIII-C Low Tensor Rank